OpenStudy (bhumi273):

a sweet contains sugar syrup up to 30% volume. 1) find approximately how much syrup would be found be fould in 45 sweets, each shaped like a spherical with diameter 2.8cm. 2) find the surface area of 1 sweet. 3) which mathematical concept is used above. 4) what is the value?

OpenStudy (bhumi273):

@Michele_Laino @Mimi_x3 @butterflydreamer @mathslover

OpenStudy (anonymous):

Find the volume of 1 sweet sphere. \(V = \dfrac{4}{3} \pi r^3\) 2.8 / 2 = 1.4, so the radius is 1.4, plug that in: \(V = \dfrac{4}{3}(3.14)(1.4^3)\) Can you simplify that?

OpenStudy (bhumi273):

yeh...

OpenStudy (michele_laino):

Hint: If V is the volume of a sweet, then total volume is 45*V

OpenStudy (bhumi273):

ok...

OpenStudy (michele_laino):

Hint: the surface area of a sphere of radius R is given by the subsequent formula: \[S = 4\;\pi \;{R^2}\]

OpenStudy (anonymous):

Okay, what do you get? @bhumi273

OpenStudy (bhumi273):

SOME... 40.7